# Conceptualization; Methodology; Resources; Investigation; Writing original draft; Writing, review, and editing

### v1

@@ 2.3 Metrics
ASD is a commonly used evaluation measure in medical image segmentation tasks. It quantifies the average distance between the surfaces of two segmented objects, typically the predicted segmentation and the ground truth.
AVD
$\max(\text{d}(A,B),\text{d}(B,A))$
@@
AVD quantifies the average difference in volume between a predicted segmentation and a reference or ground truth segmentation..
SNR
$\mathrm{10\log_{10}\Big{(}\frac{Signal}{Noise}\Big{)}_{dB}}$
[…]
SNR is a measure of the quality of the speech signal. It is commonly used to evaluate the quality of the stego-speech (the speech signal after the hidden information has been embedded). A lower SNR indicates that the steganography technique has introduced more distortion to the speech signal.
ASE
$\frac{\sum E_{ij}}{N}$
Is the average surface error (ASE). The distances between corresponding points on the measured surface and the reference surface are computed using $E_{ij}=|I_{1}(i,j)-I_{2}(k,l)|$, known as point to point error (P2PE). $N$ is the total number of correspondence points. Normalize average error by dividing by the imaging system’s dynamic range. $(i,j)$, and $(k,l)$ represent pixel values in the first and second images, respectively.
ILD
$20\log_{10}\left(\frac{L}{R}\right)$
[…]
ILD a psychoacoustic metric, measuring sound level differences between left (L) and right (R) ears, reflects cues essential for sound localization, enhancing spatial awareness in auditory perception for directional sound source identification.
MEE
$\frac{1}{N}\sum_{i=1}^{N}|L_{i}-\hat{L}_{i}|$
[…]
MEE measures the average absolute difference between the true $L_{i}$ and estimated $\hat{L}_{i}$ endpoint locations across multiple utterances $N$.
MMSE
Estimator $\hat{x}_{MMSE}=E[X|Y]$
[…]
MMSE is a statistical estimation technique used in speech enhancement to minimize the mean square error between the estimated $Y$ and true $X$ clean speech signals
STOI
$\frac{\sum_{t=1}^{T}g(t)\cdot\text{STOI}_{\text{frame}}(t)}{\sum_{t=1}^{T}g(t)}$
[…]
STOI is a metric used to assess the intelligibility of time-frequency weighted noisy speech. It is based on the idea that human speech perception relies on the availability of important acoustic features in short time frames [37].
NCM
[…]
NCM
SDR, SAR, and SIR
$10log_{10}\frac{{\left\|s_{target}\right\|}^{2}}{{\left\|e_{total}\right\|}^{2}},\hskip 8.5359pt10log_{10}\frac{{\left\|s_{target}+e_{inter}+e_{noise}\right\|}^{2}}{{\left\|e_{artifacts}\right\|}^{2}},\newline \hskip 8.5359pt10log_{10}\frac{{\left\|s_{target}\right\|}^{2}}{{\left\|e_{inter}\right\|}^{2}}$
[…]
SDR, source-to-artifact ratio (SAR), and source-to-interference ratio (SIR) are metrics objectively assess and compare speech source-separation algorithms based on accuracy and minimization of distortions and interference. source-to-distortion ratio (SDR) gauges source separation quality by comparing true source power to introduced distortion. SAR evaluates source separation from artifacts or noise, while SIR measures the ratio of true source power to interference after separation.
3 Taxonomy of CI-based AI techniques
Several artificial intelligence techniques have been employed to enhance the efficacy of CIs. While some rely on 1D data, others process information in a 2D image format. Figure 4 summarizes all AI algorithms utilized, alongside the features employed and hybrid AI methodologies. Additionally, Table 4 provides a summary of DL-based techniques utilized in CI hearing devices.
Figure 4: Taxonomy of the employed AI techniques for CI.
3.1 CI-based AI Implementation
[…]
CI programming involves adjusting device settings to optimize sound perception for individual users. This includes setting stimulation levels, electrode configurations, and signal processing parameters to enhance speech understanding and auditory experiences based on patient feedback and objective measures. In 2010, govaerts et al. [38] describes the development of an intelligent agent, called fitting to outcomes expert (FOX), for optimizing CI programming as described in Figure 6. The agent analyzes map settings and psycho-acoustic test results to recommend and execute modifications to improve outcomes. The tool focuses on an outcome-driven approach, reducing fitting time and improving the quality of fitting. It introduces principles of AI in the CI fitting process. The study proposed objective measures and group electrode settings as strategies to reduce fitting time.
@@ 3.1 CI-based AI Implementation
Similarly in [39, 40, 41, 42, 43], all have employed FOX for programming CI. Vaerenberg et al. [39] discusses the use of FOX for programming CI sound processors in new users. FOX modifies maps based on specific outcome measures using heuristic logic and deterministic rules. The study showed positive results and optimized performance after three months of programming, with good speech audiometry and loudness scaling outcomes. The paper highlights the importance of individualized programming parameters and the need for outcome-based adjustments rather than relying solely on comfort. In [40], computer-assisted CI fitting using FOX assessed its impact on speech understanding. Results from 25 recipients showed that 84% benefited from suggested map changes, significantly improving speech understanding thanks to the learning capacity of FOX. This approach offers standardized, systematic CI fitting, enhancing auditory performance.
The COCH gene, referred to as the Cochlin gene, is responsible for encoding the cochlin protein situated on chromosome 14 in humans, primarily expressed in the inner ear. Cochlin predominantly functions within the cochlea, a spiral-shaped structure involved in the process of hearing, contributing to its structural integrity and proper operation. Wathour et al. in [41] discusses the use of AI in CI fitting through two case studies. The first case involves a 75-year-old lady who received a left ear implant due to gradual and severe hearing loss in both ears without a clear cause. In the second case, a 72-year-old man with a COCH gene mutation causing profound hearing loss in both ears underwent a right ear implant, to assess whether CI programming using the AI software FOX application, could improve CI performance. The results showed that CI-assisted fitting led to improvements in auditory outcomes for adult CI recipients who had previously undergone manual fitting. The AI suggestions helped improve word recognition scores and loudness scaling curves. Likewise, Waltzman et al. [42] incorporate AI in programming CIs, aimed to assess, on fifty-five adult’s CI , the performance and standardization of AI-based programming. The results showed that the AI-based FOX for some patients performed better, while others had similar results, yet he majority preferred the FOX system.
@@
Moreover, the research in [47] focuses on imputing missing values in CI candidate audiometric data. This study assessed the performance of various imputation algorithms using a dataset of 7,451 audiograms from CI patients. The results showed that the quantity of missing data affected the imputation performance, with greater amounts leading to poorer results. The distribution of sparsity in the audiometric data was found to be non-uniform, with inter-octave frequencies being less commonly tested. The multiple imputation by chained equations (MICE) method, safely imputed up to six missing data points in an 11-frequency audiogram, consistently outperformed other models. This study highlights the importance of imputation techniques in maximizing datasets in hearing healthcare research. Xu et al. in [48] explores the objective discrimination of bimodal speech using frequency following responsess. The study investigates the neural encoding of fundamental frequency ($f_{0}$), called also pitch [49], and temporal fine structure cues (TFSC) in simulated bimodal speech conditions. The results show that increasing acoustic bandwidth enhances the neural representation of $f_{0}$ and TFSC components in the non-implanted ear. Moreover, ML algorithms successfully classify and discriminate FFRs based on spectral differences between vowels. The findings suggest that the enhancement of $f_{0}$ and TFSC neural encoding with increasing bandwidth is predictive of perceptual bimodal benefit in speech-in-noise tasks. FFRs may serve as a useful tool for objectively assessing individual variability in bimodal hearing. The research conducted by Crowson et al. [50] aimed to predict postoperative CI performance using supervised ML. The authors used neural networks and decision tree-based ensemble algorithms on a dataset of 1,604 adults who received CIs. They included 282 text and numerical variables related to demographics, audiometric data, and patient-reported outcomes. The results showed that the neural network model achieved a 1-year postoperative performance prediction root mean square error (RMSE) of 0.57 and classification accuracy of 95.4%. When both text and numerical variables were used, the RMSE was 25.0% and classification accuracy was 73.3%. The study identified influential variables such as preoperative sentence-test performance, age at surgery, and specific questionnaire responses. The findings suggest that supervised ML can predict CI performance and provide insights into factors affecting outcomes. In the same context of prediction, Mikulskis et al. [51] focuse on predicting the attachment of broad-spectrum pathogens to coating materials for biomedical devices as illustrated in Figure 6. The authors employ ML methods to generate quantitative predictions for pathogen attachment to a large library of polymers. This approach aims to accelerate the discovery of materials that resist bacterial biofilm formation, reducing the rate of infections associated with medical devices. The study highlights the need for new materials that prevent bacterial colonization and biofilm development, particularly in the context of antibiotic resistance. The results demonstrate the potential of ML in designing polymers with low pathogen attachment, offering promising candidate materials for implantable and indwelling medical devices. Similarly, Alohali et al. [52] focuses on using ML algorithms to predict the post-operative electrode impedances in CI patients. The study used a dataset of 80 pediatric patients and considered factors such as patient age and intraoperative electrode impedance. The results showed that the best algorithm varied by channel, with Bayesian linear regression and neural networks providing the best results for 75% of the channels. The accuracy level ranged between 83% and 100% in half of the channels one year after surgery. Additionally, the patient’s age alone showed good prediction results for 50% of the channels at six months or one year after surgery, suggesting it could be a predictor of electrode impedance.
@@
3.3 CNN-based methods
@@
This paper [59] focuses on advancing laser bone ablation in microsurgery using 4D optical coherence tomography (OCT). The challenge lies in automatic control without external tracking systems. The paper introduces a 2.5D scene flow estimation method using CNN for OCT images, enhancing laser ablation control. A two-stage approach involves lateral scene flow computation followed by depth flow estimation. Training is semi-supervised, combining ground truth error and reconstruction error. The method achieves a mean endpoint error (MEE) of (4.7 $\pm$ 3.5) voxel, enabling markerless tracking for image guidance and automated laser ablation control in minimally invasive cochlear implantation.
Table 4: Summary of some proposed methods based on differents DL techniques. When comparing the work with numerous existing schemes, only the best-performing one will be highlighted.
Ref.
DL method
Domain
Dataset
Explanation
CTODLM
Best performance
Improvement
[15]
FCN
Speech
MHINT
Experiment the effectiveness of FCN in restoring clean speech signals from noisy counterparts and enhancing the intelligibility of speech in EAS devices.
DDAE
STOI=0.85 (SNR=9dB)
STOI=0.03
[19]
CGRU
Speech
PhonDat 1
Using of onset and offset transitions in speech analysis and The input sequences are 3D-channel inputs
CGRU
Pre=0.68 ± $\pm\ 0.22$
@@
[59]
CNN
Image
Created
Multiple stages and follows a pyramidal network architecture inspired by SPyNet, which is designed for optical flow estimation.
MEE=4.7± $\pm\ 3.5$
[20]
multi-channel CNN
Speech
PhonDat 1
Developing a methodology for the automatic detection of speech disorders in CI users using a 2D-channel CNN.
SVM
Acc=83.4,
@@
[16]
DDAE
Speech
MHINT
The architecture of the DDAE model consisting of five layers with 500 neurons each. Objective evaluations were conducted using STOI and NCM, and a listening test (PESQ)
KLT
NCM=0.25 (SNR=9dB )
@@
[60]
Res-CA
Image
Created
Introduces Res-CA block, GCPFE, ACE loss, and DS mechanism to enhance the DAE’s performance and robustness for compressing CI stimulation patterns.
3D-DSD
DCS=89.91±$\pm 1.77$
@@
[61]
MMS [^14] https://github.com/AngeLouCN/Min_Max_Similarity
Image
Created
The classifiers extract features to create negative pairs, while projectors transform unlabeled predictions into high-dimensional features to measure pixel-wise consistency.
TransUNet
DCS=0.920
@@
[62]
MARGAN
Image
Created
MARGAN is a 3D generative adversarial network (GAN)-based approach that uses simulated training data to reduce metal artifacts in CT images.
marLI
RMSE =0.12
@@
[63]
cGAN
Image
Created
Active shape model-based method used for segmenting intra-cochlear anatomical structures.
ASE=0.173
ASE=-50%
[25]
MIMO-TasNet using RNN
Speech
WSJ0-2mix,
@@
A speech separation framework using TasNet and recurrent neural network (RNN)- eigenvector decomposition (EVD) to preserve spatial cues for CI users.
MIMO-TasNet
$\Delta$ interaural level difference (ILD)=0.38 decibel (dB)
$\Delta$ILD=-0.69 dB
[31]
DNN
Speech
iKala, MUSDB
A DNN algorithm enhances music perception for CI users, addressing pitch and timbre limitations. Objective measurements and listener experiments reveal preferences for enhanced vocals.
SDR=8dB (with iKala)
@@
[64]
Hybrid DNN
Speech
Created
Hearing training system consists of a speech training database, automatic lip-reading using a hybrid neural network, comparison of lip shapes, and providing a standard lip-reading sequence.
Acc=98 (gesture recognition) Acc=87 (lip-reading recognition)
[65]
ESCSO-based LSTM
ERP of EEG
Created
LSTM-ESCSO is a method that combines long short term memory (LSTM) network with enhanced swarm based crow search optimization (ESCSO) to improve the performance of predicting CI scores.
LSTM
Acc=96.03
Acc=+7.14
[66]
FOX
Speech
Created
The study investigates the utilization of AI in remote CI fitting, demonstrates the feasibility of AI-based fitting with remote supervision, highlighting its potential for streamlining the fitting process.
Manual fitting
66% performed better with FOX
[43]
FOX
Speech
Created
A study comparing manual and computer-assisted CI fitting.
Manual fitting
82% performed better with FOX
[67]
FOX
Speech
Created
Examined the impact of an AI-based CI programming tool on experienced CI patients. The AI-generated CI map led to improved auditory outcomes.
Manual fitting
89% preferring FOX’s map
Abbreviations: Project link (PL), Not available (NA), Compared to other DL methods (CTODLM).
3.4 GAN-based methods

### v3

@@ II-C Metrics
SDR, SAR, and SIR $10log_{10}\frac{{\left\|s_{target}\right\|}^{2}}{{\left\|e_{total}\right\|}^{2}},\hskip 8.5359pt10log_{10}\frac{{\left\|s_{target}+e_{inter}+e_{noise}\right\|}^{2}}{{\left\|e_{artifacts}\right\|}^{2}},\newline \hskip 8.5359pt10log_{10}\frac{{\left\|s_{target}\right\|}^{2}}{{\left\|e_{inter}\right\|}^{2}}$Ac SDR, source-to-artifact ratio (SAR), and source-to-interference ratio (SIR) are metrics objectively assess and compare speech source-separation algorithms based on accuracy and minimization of distortions and interference. source-to-distortion ratio (SDR) gauges source separation quality by comparing true source power to introduced distortion. SAR evaluates source separation from artifacts or noise, while SIR measures the ratio of true source power to interference after separation.
III Taxonomy of CI-based AI techniques
Several artificial intelligence techniques have been employed to enhance the efficacy of CIs. While some rely on 1D data, others process information in a 2D image format. Figure 4 summarizes all AI algorithms utilized, alongside the features employed and hybrid AI methodologies. Additionally, Table LABEL:tab provides a summary of DL-based techniques utilized in CI hearing devices.
Figure 4: Taxonomy of the employed AI techniques for CI.
III-A CI-based AI Implementation
@@
CI programming involves adjusting device settings to optimize sound perception for individual users. This includes setting stimulation levels, electrode configurations, and signal processing parameters to enhance speech understanding and auditory experiences based on patient feedback and objective measures. In 2010, Govaerts et al. [41] described the development of an intelligent agent, called fitting to outcomes expert (FOX), for optimizing CI programming, as illustrated in Figure 5. The agent analyzes map settings and psychoacoustic test results to recommend and execute modifications to improve outcomes. The tool focuses on an outcome-driven approach, reducing fitting time and improving the quality of fitting. It introduces principles of AI into the CI fitting process. The study proposed objective measures and group electrode settings as strategies to reduce fitting time.
@@
Similarly, in [42, 43, 44, 45, 46], all have employed FOX for programming CI. Vaerenberg et al. [42] discusses the use of FOX for programming CI sound processors in new users. FOX modifies maps based on specific outcome measures using heuristic logic and deterministic rules. The study showed positive results and optimized performance after three months of programming, with good speech audiometry and loudness scaling outcomes. The paper highlights the importance of individualized programming parameters and the need for outcome-based adjustments rather than relying solely on comfort. In [43], computer-assisted CI fitting using FOX assessed its impact on speech understanding. Results from 25 recipients showed that 84% benefited from suggested map changes, significantly improving speech understanding thanks to the learning capacity of FOX. This approach offers standardized, systematic CI fitting, enhancing auditory performance.
The COCH gene, referred to as the Cochlin gene, is responsible for encoding the cochlin protein situated on chromosome 14 in humans, primarily expressed in the inner ear. Cochlin predominantly functions within the cochlea, a spiral-shaped structure involved in the process of hearing, contributing to its structural integrity and proper operation. Wathour et al. in [44] discuss the use of AI in CI fitting through two case studies. The first case involves a 75-year-old lady who received a left ear implant due to gradual and severe hearing loss in both ears without a clear cause. In the second case, a 72-year-old man with a COCH gene mutation causing profound hearing loss in both ears underwent a right ear implant to assess whether CI programming using the AI software FOX application could improve CI performance. The results showed that AI-assisted fitting led to improvements in auditory outcomes for adult CI recipients who had previously undergone manual fitting. The AI suggestions helped improve word recognition scores and loudness scaling curves. Similarly, Waltzman et al. [45] incorporate AI in programming CIs, aiming to assess the performance and standardization of AI-based programming on fifty-five adult CI recipients. The results showed that the AI-based FOX system performed better for some patients, while others had similar results; however, the majority preferred the FOX system.
@@
Moreover, the research in [50] focuses on imputing missing values in CI candidate audiometric data. This study assessed the performance of various imputation algorithms using a dataset of 7,451 audiograms from CI patients. The results showed that the quantity of missing data affected the imputation performance, with greater amounts leading to poorer results. The distribution of sparsity in the audiometric data was found to be non-uniform, with inter-octave frequencies being less commonly tested. The multiple imputation by chained equations (MICE) method, safely imputed up to six missing data points in an 11-frequency audiogram, consistently outperformed other models. This study highlights the importance of imputation techniques in maximizing datasets in hearing healthcare research. Xu et al. in [51] explores the objective discrimination of bimodal speech using frequency following responsess. The study investigates the neural encoding of fundamental frequency ($f_{0}$), called also pitch [52], and temporal fine structure cues (TFSC) in simulated bimodal speech conditions. The results show that increasing acoustic bandwidth enhances the neural representation of $f_{0}$ and TFSC components in the non-implanted ear. Moreover, ML algorithms successfully classify and discriminate FFRs based on spectral differences between vowels. The findings suggest that the enhancement of $f_{0}$ and TFSC neural encoding with increasing bandwidth is predictive of perceptual bimodal benefit in speech-in-noise tasks. FFRs may serve as a useful tool for objectively assessing individual variability in bimodal hearing. The research conducted by Crowson et al. [53] aimed to predict postoperative CI performance using supervised ML. The authors used neural networks and decision tree (DT)-based ensemble algorithms on a dataset of 1,604 adults who received CIs. They included 282 text and numerical variables related to demographics, audiometric data, and patient-reported outcomes. The results showed that the neural network model achieved a 1-year postoperative performance prediction root mean square error (RMSE) of 0.57 and classification accuracy of 95.4%. When both text and numerical variables were used, the RMSE was 25.0% and classification accuracy was 73.3%. The study identified influential variables such as preoperative sentence-test performance, age at surgery, and specific questionnaire responses. The findings suggest that supervised ML can predict CI performance and provide insights into factors affecting outcomes. In the same context of prediction, Mikulskis et al. [54] focuse on predicting the attachment of broad-spectrum pathogens to coating materials for biomedical devices as illustrated in Figure 6. The authors employ ML methods to generate quantitative predictions for pathogen attachment to a large library of polymers. This approach aims to accelerate the discovery of materials that resist bacterial biofilm formation, reducing the rate of infections associated with medical devices. The study highlights the need for new materials that prevent bacterial colonization and biofilm development, particularly in the context of antibiotic resistance. The results demonstrate the potential of ML in designing polymers with low pathogen attachment, offering promising candidate materials for implantable and indwelling medical devices. Similarly, Alohali et al. [55] focuses on using ML algorithms to predict the post-operative electrode impedances in CI patients. The study used a dataset of 80 pediatric patients and considered factors such as patient age and intraoperative electrode impedance. The results showed that the best algorithm varied by channel, with Bayesian linear regression and neural networks providing the best results for 75% of the channels. The accuracy level ranged between 83% and 100% in half of the channels one year after surgery. Additionally, the patient’s age alone showed good prediction results for 50% of the channels at six months or one year after surgery, suggesting it could be a predictor of electrode impedance.
@@
Recently, Zeitler et al. [56] developed supervised ML classifiers to predict acoustic hearing preservation in patients undergoing CI surgery. The classifiers were trained using preoperative clinical data from 175 patients. The analysis revealed associations between various factors and hearing preservation outcomes. The random forest classifier demonstrated the highest mean performance in predicting outcomes. ML showed potential for predicting residual acoustic hearing preservation and improving clinical decision-making in cochlear implantation.
III-C CNN-based methods
@@
This paper [63] focuses on advancing laser bone ablation in microsurgery using 4D optical coherence tomography (OCT). The challenge lies in automatic control without external tracking systems. The paper introduces a 2.5D scene flow estimation method using CNN for OCT images, enhancing laser ablation control. A two-stage approach involves lateral scene flow computation followed by depth flow estimation. Training is semi-supervised, combining ground truth error and reconstruction error. The method achieves a mean endpoint error (MEE) of (4.7 $\pm$ 3.5) voxel, enabling markerless tracking for image guidance and automated laser ablation control in minimally invasive cochlear implantation. Recently, Almansi et al. [64] presents a radiological software prototype for detecting and classifying normal and malformed inner ear anatomy using cropping algorithms and CNN to analyze CT images. The software achieved an average accuracy of 92.25% for cropping inner ear volumes and an AUC of 0.86 for classifying normal and abnormal anatomy. Additionally, Jehn et al. [65] aimed to improve auditory attention decoding (AAD) for CI users using a CNN. EEG data from 25 CI users showed that the CNN decoder achieved a maximum decoding accuracy of 74% for a decision window of 60 seconds. Besides, the work in [66] introduces a method for detecting dysphonic voice using cochleagram images and a pre-trained CNN, achieving 95% accuracy with sentence samples. However, [67] proposes a method combining high-resolution spiral CT scanning with DL technique for diagnosing auriculotemporal and ossicle-related diseases. The study utilizes CNN-UNet model to extract sub-pixel information from medical photos of the cochlea. The results demonstrate that this approach improves diagnostic efficiency and enhances understanding of these complex diseases.
TABLE IV: Summary of some proposed methods based on different DL techniques. When comparing the work with numerous existing schemes, only the best-performing one will be highlighted.
Ref. DL method Domain Dataset Explanation CTODLM Best performance Improvement
[17] FCN Speech MHINT Experiment the effectiveness of FCN in restoring clean speech signals from noisy counterparts and enhancing the intelligibility of speech in EAS devices. DDAE STOI=0.85 (SNR=9dB) STOI=0.03
[21] CGRU Speech PhonDat 1 Using of onset and offset transitions in speech analysis and The input sequences are 3D-channel inputs CGRU Pre=0.68 ± $\pm\ 0.22$
@@
[63] CNN Image Created Multiple stages and follows a pyramidal network architecture inspired by SPyNet, which is designed for optical flow estimation. MEE=4.7± $\pm\ 3.5$
[22] multi-channel CNN Speech PhonDat 1 Developing a methodology for the automatic detection of speech disorders in CI users using a 2D-channel CNN. SVM Acc=83.4,
@@
[18] DDAE Speech MHINT The architecture of the DDAE model consisting of five layers with 500 neurons each. Objective evaluations were conducted using STOI and NCM, and a listening test (PESQ) KLT NCM=0.25 (SNR=9dB )
@@
[68] Res-CA Image Created Introduces Res-CA block, GCPFE, ACE loss, and DS mechanism to enhance the DAE’s performance and robustness for compressing CI stimulation patterns. 3D-DSD DCS=89.91±$\pm 1.77$
@@
[69] MMS [^16] https://github.com/AngeLouCN/Min_Max_Similarity Image Created The classifiers extract features to create negative pairs, while projectors transform unlabeled predictions into high-dimensional features to measure pixel-wise consistency. TransUNet DCS=0.920
@@
[70] MARGAN Image Created MARGAN is a 3D generative adversarial network (GAN)-based approach that uses simulated training data to reduce metal artifacts in CT images. marLI RMSE =0.12
@@
[71] cGAN Image Created Active shape model-based method used for segmenting intra-cochlear anatomical structures. ASE=0.173 ASE=-50%
[27] MIMO-TasNet using RNN Speech WSJ0-2mix,
@@
A speech separation framework using TasNet and recurrent neural network (RNN)- eigenvector decomposition (EVD) to preserve spatial cues for CI users. MIMO-TasNet $\Delta$ interaural level difference (ILD)=0.38 decibel (dB) $\Delta$ILD=-0.69 dB
[33] DNN Speech iKala, MUSDB Enhances music perception for CI users, addressing pitch and timbre limitations. Objective measurements and listener experiments reveal preferences for enhanced vocals. SDR=8dB (with iKala)
@@
[72] Hybrid DNN Speech Created Hearing training system consists of a speech training database, automatic lip-reading using a hybrid neural network, comparison of lip shapes, and providing a standard lip-reading sequence. Acc=98 (gesture recognition) Acc=87 (lip-reading recognition)
[73] ESCSO-based LSTM ERP of EEG Created Method that combines long short term memory (LSTM) network with enhanced swarm based crow search optimization (ESCSO) to improve the performance of predicting CI scores. LSTM Acc=96.03 Acc=+7.14
[74] FOX Speech Created The study investigates the potential of utilization of AI in remote CI fitting, demonstrates the feasibility of AI-based fitting with remote supervision. Manual fitting 66% performed better with FOX
[46] FOX Speech Created A study comparing manual and computer-assisted CI fitting. Manual fitting 82% performed better with FOX
[75] FOX Speech Created Examined the impact of an AI-based CI programming tool on experienced CI patients. The AI-generated CI map led to improved auditory outcomes. Manual fitting 89% preferring FOX’s map
[76] DNN and
self-attention
Stimulus Created Using self-attention modules, speech reconstruction from neural activity improved by leveraging long-span neural data, outperforming CNN and multilayer perceptrons (MLP) models, enhancing hearing evaluation and neurosurgery. CNN and MLP ESTOI=0.239 ESTOI=+0.009
[77] Bi-LSTM Audiograms Created A bidirectional long short term memory (Bi-LSTM) classifies tonal audiometry data, aiding general practitioners and reducing audiologists’ burden by diagnosing hearing loss types independently. LSTM Acc=99.33% Acc=+1.04%
[65] CNN and SVM Speech Created A CNN significantly improves AAD for CI users, enhancing listening experiences by steering auditory prostheses towards target speech. Acc=74%
[…]
III-D GAN-based methods